OpenStudy (anonymous):

Find the derivative of f(x) = -11/x at x = 9.

OpenStudy (anonymous):

Can you use the power rule to get the derivative of f(x)?

OpenStudy (anonymous):

How would I do that?

OpenStudy (cwrw238):

-11/x = -11x^-1 now use the power rule

OpenStudy (anonymous):

You can rewrite f(x) = -11x^(-1), is this a more recognisable form?

OpenStudy (anonymous):

yes, thank you @aleroth .. Okay, so... -1 \[-1 \times -11x ^{-2}\]

OpenStudy (anonymous):

That you have now is the derivative of f(x), aka f'(x). The question wants you to plug in 9 into it as well to get the "derivative when x = 9"

OpenStudy (anonymous):

so just plug in 9 for x? that's it?

OpenStudy (anonymous):

Yep. Also just a note; f'(x) can also be written as 11/x^2. Know that 1/x = x^-1.

OpenStudy (anonymous):

so the final answer should be 11/81? correct?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thank you so much!!(:

OpenStudy (anonymous):

Glad to help. :)